OpenStudy (anonymous):

a firefighter jumps out the window into a net.if the window is 25m above the net, a) at what velocity does the fighter hit the net? b) how long will it take the firefighter to hit the west?

OpenStudy (anonymous):

well, the acceleration would be -9.8 m/s^2 do to gravity.

OpenStudy (anonymous):

what is the given ? i need formula? thank u for helping me ...

OpenStudy (anonymous):

please help me ???? i need the equation ??? thank u

OpenStudy (anonymous):

1) y=4(x+1)*2 ,2) y=-5(x-2)*2+15, 3)y=-2(+2)*2-1, 4) y=3(x-6)*2-11,5) y=(x-10)*2-25, transform the ff functions in the form y=ax*2+bx+c

OpenStudy (anonymous):

v^2=u^2+2*a*s u=initial vel=0 v final vel a =g s=25 v=sqrt2*9.8*25 v=u+a*t u=0 a=g t=v/g

OpenStudy (anonymous):

i dont understand can u explain to me ??? sorry

OpenStudy (anonymous):

You probably want to take this to the physics section.

OpenStudy (anonymous):

please more explanation in physics please.....

OpenStudy (anonymous):

Vf^2 = Vi^2 + 2ad ; final velocity = Vf, initial velocity is Vi, a = acceleration, d = displacement Because he is just free falling out of a window, his Vi is 0 His displacement is 25 m because he's 25m about the net. Acceleration is constant -9.8 m/s^2 Plug in the components to figure out Vf. Then using Vf = Vi + at plug in Vf, Vi, and a to figure out how long it takes.